OpenStudy (anonymous):

(7th grade math) Only adding fractions... I just need an answer checked...or question idk A painter used 5/6 gallon of paint for one job and 1/4 gallon for another job. How many gallons of paint did the painter use for both jobs? A. 1/2 gal B. 3/5 gal C. 1 1/12 gal D. 1 1/6 gal I got 1 1/12

OpenStudy (whpalmer4):

\[\frac{5}{6} + \frac{1}{4}\]looks like \(12\) would be a convenient common denominator: \(2*6 = 12\) and \(3*4 = 12\) \[\frac{2}{2}*\frac{5}{6} + \frac{3}{3}*\frac{1}{4} = \frac{10}{12} + \frac{3}{12} = \frac{ 13} { 12} = 1\frac{1}{12}\checkmark\]

OpenStudy (anonymous):

Thanks @whpalmer4 ^_^

OpenStudy (whpalmer4):

Another way of doing this: \[\frac{5}{6}+\frac{1}{4} = x\]\[4*6*\frac{5}{6} + 4*6*\frac{1}{4} = 4*6*x\]\[20 + 6 = 24x\]\[x = \frac{26}{24} = \frac{13}{12} = 1\frac{1}{12}\]